OpenStudy (anonymous):

there are 19 parts bran flakes and 1 part raisins in a certain cereal. if the volume of the box is 32oz,how many ounces of each food are in the box?

OpenStudy (anonymous):

19x + x = 32 20 x = 32 x = 32/20 bran flakes: 32*19/20 raisins is 32/20 Simplify the fractions

OpenStudy (anonymous):

19:1 32 ounces / (19+1) = 1.6 ounces per unit. Bran = 1.6 * 19 ounces Raisins = 1.6 * 1 ounces